OpenStudy (anonymous):

If 1/a+1/b=1/c and ab=c the what is the average of a and b?

OpenStudy (anonymous):

If by average u mean arithmetic mean then it'd be 1/2

OpenStudy (anonymous):

Please elaborate the method you used

OpenStudy (anonymous):

Explanation: \[\frac{ 1 }{ a }+\frac{ 1 }{ b }= \frac{ 1 }{ c }\] this means, \[\frac{ a+b }{ a*b }=\frac{ 1 }{ c }\] this implies \[ac+bc=ab\] substituting this value of ab in the second equation, ab = c, we get, \[a ^{2}b+ab ^{2}=ab\] dividing this entire equation by ab, we get, \[a+b=1\] this is the sum of a and b, the average should be sum of the values / no. of values, therefore, the average is \[\frac{ Sum of values }{ no. of values }=\frac{ a+b }{ 2 }=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Thank you! Best Explanation

OpenStudy (anonymous):

You are welcome!

OpenStudy (anonymous):

hi mruna_cp, could you please explain how you got \[a ^{2}b + ab ^{2}=ab\]? Thank you so much